INAF queue observing, 2015 Jan 12-13

Observer: E. Sani, S. Bisogni
Telescope Operator: G. Bechetti
Support Astronomer: Dave Thompson (on call)
Instrument: LUCI

Summary

At the beginning of the night the dome stays closed due to humidity.

Data Summary

Project_RUN Objects Filter ExpTime Notes
LUCI Calibrations

Details

07:38 we start running LUCI calibration (arcs)
  • executeLUCIScript.sh A1_LS_Ks.arc luci.0001-0004
  • executeLUCIScript.sh Target_11E.arc luci.0005-0010
  • executeLUCIScript.sh Target_11F.arc luci.0011-0014

the counts are low; we repeat arc calibrations with longer exposure times

  • executeLUCIScript.sh A1_LS_Ks.arc luci.0015-0018
  • executeLUCIScript.sh Target_11E.arc luci.00019-0024
  • executeLUCIScript.sh Target_11D.arc luci.0025-0028

in the last arc counts are still low, we repeat Target_11D.arc with longer exposure time
  • executeLUCIScript.sh Target_11D.arc luci.0029-0032 some lines are saturated

-- %USERSIG{SusannaBisogni - 2015-01-13}%

Comments

 
Topic revision: r4 - 19 May 2015, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback